OpenStudy (anonymous):

a, b, c, 63....... In the sequence above, each term after the first is 1 less than the square of the term preceding it. What is the value of a?

OpenStudy (anonymous):

c^2 =63+1

OpenStudy (anonymous):

b^2 = c+1

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

thanks:))

OpenStudy (anonymous):

2